• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name R. Michaels

    Log entry time 17:54:12 on March 07, 2010

    Entry number 311013

    keyword=HAPPEX DAQ status

    Since Friday Luis and I did the following to complete the HAPPEX DAQ.
    There is still a mystery (# 2).

    1. Installed all ADCs and tested, halog 310932

    2. Chased the intermittent "bad" data in ADC18 in CH crate.
    See halog 310932,   310933,     310962,   310993  
    Finally "solved" it with a dirty trick.   Need to brainstorm with
    DAQ group and do more tests this week, as time permits.

    3. Found Lumi cables. We should get Zafar to check Lumi's. (HV on/off test.)

    4. I would like to thoroughly check the helicity-related signals, too.   Zafar has started.

    5. I'm not sure what the status of the noise studies is.   May need to do more.

    6. Put main DAQ configs on adaql5.   halog 311011

    7. Protocol for using DAQ:
    a) Make advance reservation if possible.   E.g. I know about the injector work Monday.
    b) Message by e-mail and halog when things are in use, or changed.


    8. DAQ basically works   except for mystery in #2.